Can you please help with this question: Two consecutive whole numbers (x and x+1) when the product of these numbers are subtracted from 47 times the first number (x), the answer is 360. what is x?
i have tried using the quadratic equation. I cannot get the answer to a whole number. my equation used to insert in the quadratic was: x(x+1)-47x-360=0.
can you please help and include all your workings so i can learn for next time. Thank you.

Answer by Alan3354(69443)

47x - x*(x+1) = 360
47x - x^2 - x - 360 = 0
x^2 - 46x + 360 = 0
(x-10)*(x-36) = 0
x = 10, 36